A train of mass 600 metric ton is pulled by a locomotive of mass 150 metric ton. The speed of train is 54 kmph. The locomotive pulls the train on level track. The force of friction on locomotive and the train is10 Newton per metric ton. Calculate the Power of the locomotive

Grade:11

1 Answers

murali
4 years ago
Frictional  force per metric ton is 10N
Frictional force on train =10×600=6000N
Froctional force on Loco =10×150=1500N
Net F =6000+1500=7500N
Power P=FV
               =7500×15=112.5Kw